• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gaskell

    Log entry time 10:32:53 on October 26, 2010

    Entry number 336801

    keyword=Status and plan for the day

    
    We completed the 2nd iteration of elastic settings last night on OWL
    shift. This morning beam went away for 8 hours of beam study. We are
    presently in Controlled Access for kinematics change and other work.
    
    Today, we will go back to DVCS production setting "Kin 3, low". This means:
    
    Theta_calo = 14.98 degrees
    Calo distance from target = 1.1 m
    
    P_HRSL = 1.494 GeV
    Theta_HRSL=14.78 degrees.
    
    Charles and I moved the calorimeter already. There is work on the HRS
    cryo system and bogies ongoing right now, so we will need to wait until
    that is complete before we set the HRS.